A swimming pool in a club is maintained sufficiently alkaline so that the

CO2 absorbed from the atmosphere produced in the pool a solution which was 2 x 10-4 M in CO3 2- after about a day of refilling the swimming. If the pool water was originally 4 x 10-3 M in Mg2+, 6 x 10-4 M in Ca2+ and 8 x 10-7 M in Fe2+, then which of the following precipitates should be formed immediately a day after refilling of the swimming pool:

(a) only MgCO3

(b) only CaCO3

(c) only FeCO3

(d) only CaCO3 and FeCO3

(e) MgCO3, CaCO3 and FeCO3


1
Expert's answer
2022-03-07T17:01:04-0500

Correct option is D)

From given values of concentations, the ionic product is calculated and compared with solubility product values.For CaCO3​,, the ionic product is [Ca2+][CO32−​]=6×10−4×2×10−4=1.2×10−9.The ionic product is less than the solubility product which is 2.8×10−9. Hence, CaCO3​, will not precipitate.For MgCO3​, the ionic product is [Mg2+][CO32−​]=4×10−3×2×10−4=8×10−7.The ionic product is more than the solubility product which is 3.5×10−8. Hence, MgCO3​ will precipitate.For FeCO3​, the ionic product is [Fe2+][CO32−​]=8×10−7×2×10−4=1.6×10−12.The ionic product is more than the solubility product which is 3.2×10−11. Hence, FeCO3​ will precipitate.

Correct answer is D